# Super resonance: Breaking the bandwidth limit of resonant modes and its application to flow control

### v1

@@ INTRODUCTION
The super resonance phenomenon can be directly observed from the frequency response function (FRF) of the version of the PSub with three coiling cycles, as shown in the amplitude and phase plots of Figs. 2b and 2c, respectively. This can be examined in reference to the lowest-frequency resonant peak shown in the figure, which is at 278 Hz. This will be considered our target resonance. Compared to the reference uncoiled case, the 3-cycle coiling has shifted the first anti-resonant trough following this resonance from 555 Hz to 826 Hz, thus increasing the out-of-phase band from $\Delta f_{\text{OoP}}=277$ Hz to $\Delta f_{\text{OoP}}=248$ Hz, i.e., almost doubling it. Thus, the coiling and multiple connectivity operation, as outlined above, has transformed this target resonance into a super resonance  [^4] It should be noted that without the described multiple structural connectivity to the flow, the coiled PSub will yield an identical FRF as its uncoiled counterpart. The sole purpose of the coiling is primarily to enable the geometric convergence of the multiple junctions to a specific point in space, namely the point that interfaces with the flow. The broadening of the out-of-phase region in the FRF is enabled by the superposition of the transfer functions.. Upon further examination, we observe that the new 826-Hz anti-resonance trough coincides with the second resonance peak appearing in the original uncoiled structure’s FRF, at practically the same frequency. Thus, the out-of-phase region associated with the 278-Hz target resonance has effectively extended to the next anti-resonance of the coiled PSub, which is at 1737 Hz—yielding an out-of-phase band of $\Delta f_{\text{OoP}}=1459$ Hz. We refer to this further widened band characteristic as a quasi-super resonance, since it practically exhibits the super-resonance behavior yet stretches its out-of-phase bandwidth further. Compared to the original configuration, the quasi super-resonant out-of-phase band is more than five times larger.
@@
Figure 4: Coiled PSub behavior by FIK integral analysis: Evaluation of various key flow parameters when the super-resonant 3-cycle coiled PSub is installed versus the rigid-wall case. (a) Streamwise distribution of the averaged skin-friction coefficient $C_{f}$, (b) streamwise evolution of the “turbulent enhancement” to $C_{f}$ computed by the FIK equation, (c) streamwise variation of the wall-normal integral of the production rate $P_{\text{r}}$, and (d) wall-normal variation of $P_{\text{r}}$ at streamwise positions corresponding to the leading (Station 1) and trailing (Station 2) edges of the control surface.
@@ CONCLUSION
In conclusion, this work presents the concept of super resonance—a new fundamental regime in resonance physics characterized by the persistence of out-of-phase modal response across a broad spectral range, despite the presence of adjacent resonant modes in the nominal configuration. In this regime, the narrowband limitation of a target mode is lifted, and the associated phase-flipping behavior with nearby modes is effectively prevented. This unique behavior is realized by a mechanism of multiple structural energy pathways spatially converging to a single point where the response is measured. This conceptual advance opens new directions in wave-matter interaction and phase-based control strategies across physics, materials science, and engineering.
Within this broader framework, we demonstrate a super-resonant PSub designed to significantly widen the frequency range over which TS instabilities are attenuated within a laminar channel flow. This entails the generation of responsive vibratory motion, at a single point interfacing with the flow, whereby the phase is contiguously negative over a frequency range that exceeds the characteristic bandwidth of a conventional structural resonance. The PSub is formed from a coiled configuration with periodic rotational locking. By enabling multiple interaction points along the coiled structure to influence that single location at the flow interface, we achieve passive stabilization across a frequency range more than five times broader than that is attainable with a corresponding uncoiled PSub. Across this entire range, the rate of production of the perturbation kinetic energy in the flow is suppressed. This broadened control window is shown to be sufficiently wide to fully cover, for a given Reynolds number, the unstable frequency band predicted from linear stability theory. This approach thus greatly enhances the effectiveness of PSubs by enabling broadband flow control—a key trait lacking across existing passive and active flow-instability control technologies [6]. As a bonus, the total height of the structure beneath the surface is reduced, which is critical for practical implementation. Upon integration with PSub concepts for downstream control [32, 33, 34], super-resonant coiled PSubs will provide a robust and tunable mechanism for delaying transition to turbulence in real-world channel and boundary layer flows. Most significantly, the concept offers a pathway towards controlling fully developed turbulent flows, whose broadband spectrum has long resisted effective control. Reaching this target will be accelerated by further enriching the design of the constituent components of super-resonant PSubs, drawing on the extensive body of knowledge in phononic crystals and elastic metamaterials [23, 24, 25, 26], including the emerging themes of topological [59], nonlocal [60], and nonlinear [61] phononics, to enable additional broadening of the out-of-phase frequency range.
@@ Acknowledgments
Figure A4 shows the relative error associated with the FIK identity along the streamwise direction for the case involving multiple TS waves (see main article). This relative error is defined as

### v4

@@ INTRODUCTION
The super resonance phenomenon can be directly observed from the frequency response function (FRF) of the version of the PSub with three coiling cycles, as shown in the amplitude and phase plots of Figs. 2b and 2c, respectively. This can be examined in reference to the lowest-frequency resonant peak shown in the figure, which is at 278 Hz. This will be considered our target resonance. Compared to the reference uncoiled case, the 3-cycle coiling has shifted the first anti-resonant trough following this resonance from 555 Hz to 826 Hz, thus increasing the out-of-phase band from $\Delta f_{\text{OoP}}=277$ Hz to $\Delta f_{\text{OoP}}=548$ Hz, i.e., almost doubling it. Thus, the coiling and multiple connectivity operation, as outlined above, has transformed this target resonance into a super resonance  [^5] It should be noted that without the described multiple structural connectivity to the flow, the coiled PSub will yield an identical FRF as its uncoiled counterpart. The sole purpose of the coiling is primarily to enable the geometric convergence of the multiple junctions to a specific point in space, namely the point that interfaces with the flow. The broadening of the out-of-phase region in the FRF is enabled by the superposition of the transfer functions.. Upon further examination, we observe that the new 826-Hz anti-resonance trough coincides with the second resonance peak appearing in the original uncoiled structure’s FRF, at practically the same frequency. Thus, the out-of-phase region associated with the 278-Hz target resonance has effectively extended to the next anti-resonance of the coiled PSub, which is at 1737 Hz—yielding an out-of-phase band of $\Delta f_{\text{OoP}}=1459$ Hz. We refer to this further widened band characteristic as a quasi-super resonance, since it practically exhibits the super-resonance behavior yet stretches its out-of-phase bandwidth further. Compared to the original configuration, the quasi super-resonant out-of-phase band is more than five times larger.
@@
Figure 4: Super-resonant coiled PSub behavior by FIK integral analysis: Evaluation of various key flow parameters when the super-resonant 3-cycle coiled PSub is installed versus the rigid-wall case. (a) Streamwise distribution of the averaged skin-friction coefficient $C_{f}$, (b) streamwise evolution of the “turbulent enhancement” to $C_{f}$ computed by the FIK equation, (c) streamwise variation of the wall-normal integral of the production rate $P_{\text{r}}$, and (d) wall-normal variation of $P_{\text{r}}$ at streamwise positions corresponding to the leading (Station 1) and trailing (Station 2) edges of the control surface.
@@ CONCLUSION
In conclusion, this work presents the concept of super resonance—a new fundamental regime in resonance physics characterized by the persistence of out-of-phase modal response across a broad spectral range, despite the presence of adjacent resonant modes in the nominal configuration. In this regime, the narrowband limitation of a target mode is effectively lifted, and the associated phase-flipping behavior with nearby modes is prevented. This unique behavior is realized by a mechanism of multiple structural energy pathways spatially converging to a single narrow spatial region where the response is measured. This conceptual advance opens new directions in wave-matter interaction and phase-based control strategies across physics, materials science, and engineering.
Within this broader framework, we demonstrate a super-resonant PSub designed to significantly widen the frequency range over which TS instabilities are attenuated within a laminar channel flow. This entails the generation of responsive vibratory motion, at an effectively single location interfacing with the flow, whereby the phase is contiguously negative over a frequency range that exceeds the characteristic bandwidth of a conventional structural resonance. The PSub is formed from a coiled configuration with periodic rotational locking. By enabling multiple interaction points along the coiled structure to influence that single location at the flow interface, we achieve passive stabilization across a frequency range more than five times broader than that is attainable with a corresponding uncoiled PSub. Across this entire range, the rate of production of the perturbation kinetic energy in the flow is suppressed. This broadened control window is shown to be sufficiently wide to fully cover, for a given Reynolds number, the unstable frequency band predicted from linear stability theory. This approach thus greatly enhances the utility of PSubs by enabling broadband flow control—a key trait lacking across existing passive and active flow-instability control technologies [6]. As a bonus, the total height of the structure beneath the surface is reduced, which is critical for practical implementation. Upon integration with PSub concepts for downstream control [32, 33, 34], super-resonant coiled PSubs will provide a robust and tunable mechanism for delaying transition to turbulence in real-world channel and boundary layer flows. Most significantly, the concept offers a pathway towards controlling fully developed turbulent flows, whose broadband spectrum has long resisted effective control. Reaching this target will be accelerated by further enriching the design of the constituent components of super-resonant PSubs, drawing on the extensive body of knowledge in phononic crystals and elastic metamaterials [23, 24, 25, 26], including the emerging themes of topological [61], nonlocal [62], and nonlinear [63] phononics, to enable additional broadening of the out-of-phase frequency range.
@@ Acknowledgments
Figure A4 shows the relative error associated with the FIK identity along the streamwise direction for the case involving multiple TS waves controlled by the 3-cycle coiled PSub (see main article). This relative error is defined as